OpenStudy (anonymous):

please help mw with this homework...due tomorrow 2/3g=-4 1/2

OpenStudy (anonymous):

\[2/3g=-4\frac{ 1 }{ 2 }\] this one?

OpenStudy (anonymous):

\[\frac23 g=-4\frac12\\\frac23g=-\frac82\times \frac12 \\g= -\frac82\times \frac12 \times \frac32\]

OpenStudy (anonymous):

@miteshchvm Hi, it's been long time since we're in the same post :D

OpenStudy (anonymous):

ya i guess so :D

OpenStudy (anonymous):

23g=−41223g=−82×12g=−82×12×32